# WOMD-Reasoning: A Large-Scale Dataset and Benchmark for Interaction and Intention Reasoning in Driving

### v2

@@ 7 Acknowledgements
This work was supported in part by Berkeley DeepDrive, NSF Grant. 2235013, and the Microsoft Accelerate Foundation Models Research Initiative.
@@ Appendix A Appendix
Then, for Motion-LLaVA, we show the prompt for obtaining the GPT Score used for evaluation (A.3). We further provide qualitative comparison of benchmarked multi-modal models listed in Figure 5 (A.4). We also provide an example of aggregated context (A.5) as well as factual description context (A.6) used in Chain-of-Thought inference ablations for reference. Finally, an ablation on details in multi-modal model fine-tuning strategy is attached (A.7).
@@ A.1 Details in Building WOMD-Reasoning
In this appendix, we show the details of how we build WOMD-Reasoning dataset, including detailed explanations on each step we take, the example of the translated motion scenarios, and the GPT prompts. The full version of codes and prompts used herein can be found at https://github.com/yhli123/WOMD-Reasoning.
@@ A.1.1 Automatic Translation of Driving Scenarios into Language Descriptions
Descriptions on map environment: The ego agent is heading towards intersection. The intersection center is 18.0 meters in front of the ego agent, and is 17.0 meters on the left of the ego agent. The intersection is a 4 way intersection. Descriptions on ego agent: The ego agent is on the 1 lane from the right, out of 3 lanes. Its current speed is 8 m/s. It is accelerating. Traffic Light for the ego agent is red. The ego agent is approaching a crosswalk 3 meters ahead. Descriptions on surrounding (other) agents: Surrounding agent # 0 is a vehicle. It is 23 meters in front of the ego agent, and is 17 meters on the left of the ego agent. It is heading right of the ego agent. Its current speed is 6 m/s. It is accelerating. It is in the intersection. It is 29 meters away from the ego agent. It is approaching a crosswalk 3 meters ahead. Surrounding agent # 2 is a vehicle. It is 11 meters on the right of the ego agent, and is 0 meters behind the ego agent. It is heading left of the ego agent. It is not moving. It is on the same side of the intersection as the ego agent. It is 34 meters away from the intersection center. $\cdots$ Descriptions on agents after 3 seconds: The following is the description of the ego and surrounding agents after 3.0 seconds: Surrounding agent # 0 will be 8 meters in front of the ego agent, and will be 5 meters on the left of the ego agent. It will be heading right of the ego agent. Its speed will be 7 m/s. It will be departing from the intersection. Looking from the agent’s current angle, it will be on the same side of the intersection. It will be 18 meters away from the intersection center. The ego agent will be 13 meters in front of the current place, and will be 2 meters on the right of the current place. It will be heading in the same direction as the current moment. Its speed will be 3 m/s. It will be departing from the intersection. Looking from the agent’s current angle, it will be on the same side of the intersection. It will be 20 meters away from the intersection center. Surrounding agent # 2 will be 13 meters on the right of the ego agent, and will be 9 meters behind the ego agent. It will be heading left of the ego agent. It will not be moving. Looking from the agent’s current angle, it will be on the same side of the intersection. It will be 34 meters away from the intersection center. $\cdots$
@@ A.1.2 GPT-based Interaction Analysis
With the language translation of WOMD as input, we further design a set of prompts to utilize ChatGPT-4 [34] to build the Q&A dataset, taking advantage of its abilities in analyzing interactions. The prompts come with 4 main parts. (1) System Prompt, which lets the GPT know its role and the format of its input, i.e. the language translation. (2) Responsibility Prompt, which describes the responsibilities of the GPT, including the questions it should ask and answer, as well as the output format it should use. This will be talked about in detail in the next paragraph. (3) Rules Prompt, which contains global rules to guide the output and the analysis. (4) In-context Prompt, which provides a few human-written examples of input-output pairs to perform in-context learning. An example of each part of the prompts can be found in Table 9, 10, 11, and 12 respectively.
The questions in the responsibility part of the prompt decide what would be present in the generated language Q&A dataset. To provide a comprehensive language dataset, and to present interaction analysis as accurately as possible, we organize the questions into a chain to enable a well-designed Chain-of-Thought [55, 55, 8]: First, we ask questions about the map environment, the ego and other agents’ motion. These questions would let the GPT provide Q&As useful for fine-tuning scene description agents while enhancing the GPT’s understanding of the input information at the same time. Then, we ask questions about the interactions that occur in the observable future period. To include traffic rule-induced and intention-induced interactions, we ask the GPT to think about the interaction between each pair of agents by answering a sequence of questions: Q1: Are the two agents vehicles, and are they close to each other with traffic rules or signals governing their movements? If the answer is yes, we request that GPT analyze their yielding relations according to the traffic rules. Otherwise, we ask Q2: Are the two agents vehicle and pedestrian respectively with intention conflicts? If so, the vehicle generally should yield to the pedestrian for safety. If both questions do not help find the interaction relations, we come to Q3: Does the scene show patterns of certain intention-induced interaction? We provide descriptions of a few common interactions like overtaking, following, etc. GPT can choose the best fit from these patterns to try to cover more interaction types. Based on this pipeline, we can provide interaction analysis with rich information on traffic rule-induced and human intention-induced interactions. After that, we can finally wrap up the interactions between the ego and all other agents to provide the intention of the ego agent.
@@ A.2 Examples of Simulated Vision Modality
Figure 7 provides some whole examples of the simulated scenarios generated by MetaDrive [24] simulator using ScenarioNet [25] traffic scenario simulation platform.
A.3 Prompts for evaluating GPT Score
We employ the OpenAI GPT-4-turbo model [34] to evaluate the performance of all models benchmarked on the WOMD-Reasoning. The GPT Score prompt is presented in Table 13. Notably, we design our GPT Score prompt using a similar template to the one from [7].
@@ A.3 Prompts for evaluating GPT Score
Table 14: Refinement of Motion-LLaVA training strategy.
LLaVA 0.821 0.678 0.484 5.82 0.771 2.3% 12.00 6.17
$\bm{-}$ Projector Alignment 0.821 0.697 0.492 5.96 0.782 2.3% 8.06 6.49
A.4 Qualitative Comparison of Benchmarked Multi-modal Models on Answering Questions in WOMD-Reasoning
@@
A.5 Example Aggregated Context for Chain-of-Thought Reasoning
@@
A.6 Example Factual Description Context for Chain-of-Thought Reasoning
@@
A.7 Ablation on Motion-LLaVA Training Strategy
The original LLaVA [28] training pipeline consists of two stages: a projector alignment stage (also known as pre-training) to align vision and language features, followed by a fine-tuning stage to enhance instruction-following capabilities. Throughout this process, the vision encoder remains frozen. Since our Motion-LLaVA builds on the architecture of the original LLaVA [28], it is intuitive to adopt a similar training approach for the WOMD-Reasoning task.
However, based on extensive experiments with fine-tuning strategies in Table 14, we find that incorporating projector alignment stage or keeping the motion vector encoder frozen throughout, which are practices employed in LLaVA [28], resulted in degraded performance on WOMD-Reasoning. Consequently, our Motion-LLaVA adopts a single-stage fine-tuning approach, where the motion vector encoder is unfrozen.

### v3

@@ 8 Acknowledgements
This work was supported in part by Berkeley DeepDrive, NSF Grant. 2235013, NSF Grant. 2235012, and the Microsoft Accelerate Foundation Models Research Initiative.
9 Impact Statements
Our work is of broad interest to the natural language processing (NLP) and autonomous driving communities. The proposed approach has no new ethical or social issues on its own, except those inherited from NLP or autonomous driving.
@@ Appendix A Appendix
Then, for Motion-LLaVA, whose language output enhances trajectory prediction as shown in Section 6, we provide further implementation details. Specifically, in A.3, we describe how agent information from driving scenarios is incorporated into the input of Motion-LLaVA. The training hyperparameters used for the downstream trajectory prediction module are presented in A.4. The prompt used to obtain the GPT Score for evaluation is provided in A.5.We further provide a qualitative comparison in A.6 of the benchmarked multi-modal models listed in Table 4, and introduce in A.7 additional multi-modal models not included in the table. Examples of aggregated context (A.8) and factual description context (A.9) used in Chain-of-Thought interaction reasoning are also provided for reference. An ablation study on the fine-tuning strategy used in Motion-LLaVA is included in A.10, along with an analysis of the data efficiency of WOMD-Reasoning in helping Motion-LLaVA learn driving-related reasoning in A.11.
@@ A.1 Details in Building WOMD-Reasoning
In this appendix, we show the details of how we build WOMD-Reasoning dataset, including detailed explanations on each step we take, the example of the translated motion scenarios, and the GPT prompts. The full version of codes and prompts used herein can be found in the supplemental materials.
@@ A.1.1 Automatic Translation of Driving Scenarios into Language Descriptions
Descriptions on map environment:
The ego agent is heading towards intersection. The intersection center is 18.0 meters in front of the ego agent, and is 17.0 meters on the left of the ego agent. The intersection is a 4 way intersection. Descriptions on ego agent:
The ego agent is on the 1 lane from the right, out of 3 lanes. Its current speed is 8 m/s. It is accelerating. Traffic Light for the ego agent is red. The ego agent is approaching a crosswalk 3 meters ahead. Descriptions on surrounding (other) agents:
Surrounding agent # 0 is a vehicle. It is 23 meters in front of the ego agent, and is 17 meters on the left of the ego agent. It is heading right of the ego agent. Its current speed is 6 m/s. It is accelerating. It is in the intersection. It is 29 meters away from the ego agent. It is approaching a crosswalk 3 meters ahead. Surrounding agent # 2 is a vehicle. It is 11 meters on the right of the ego agent, and is 0 meters behind the ego agent. It is heading left of the ego agent. It is not moving. It is on the same side of the intersection as the ego agent. It is 34 meters away from the intersection center.
$\cdots$
Descriptions on agents after 3 seconds:
The following is the description of the ego and surrounding agents after 3.0 seconds: Surrounding agent # 0 will be 8 meters in front of the ego agent, and will be 5 meters on the left of the ego agent. It will be heading right of the ego agent. Its speed will be 7 m/s. It will be departing from the intersection. Looking from the agent’s current angle, it will be on the same side of the intersection. It will be 18 meters away from the intersection center. The ego agent will be 13 meters in front of the current place, and will be 2 meters on the right of the current place. It will be heading in the same direction as the current moment. Its speed will be 3 m/s. It will be departing from the intersection. Looking from the agent’s current angle, it will be on the same side of the intersection. It will be 20 meters away from the intersection center. Surrounding agent # 2 will be 13 meters on the right of the ego agent, and will be 9 meters behind the ego agent. It will be heading left of the ego agent. It will not be moving. Looking from the agent’s current angle, it will be on the same side of the intersection. It will be 34 meters away from the intersection center.
$\cdots$
@@ A.1.2 GPT-based Interaction Analysis
With the language translation of WOMD as input, we further design a set of prompts to utilize ChatGPT-4 (OpenAI et al., 2024) to build the Q&A dataset, taking advantage of its abilities in analyzing interactions. The prompts come with 4 main parts. (1) System Prompt, which lets the GPT know its role and the format of its input, i.e. the language translation. (2) Responsibility Prompt, which describes the responsibilities of the GPT, including the questions it should ask and answer, as well as the output format it should use. This will be talked about in detail in the next paragraph. (3) Rules Prompt, which contains global rules to guide the output and the analysis. (4) In-context Prompt, which provides a few human-written examples of input-output pairs to perform in-context learning. An example of each part of the prompts can be found in Table 8, 9, 10, and 11 respectively.
The questions in the responsibility part of the prompt decide what would be present in the generated language Q&A dataset. To provide a comprehensive language dataset, and to present interaction analysis as accurately as possible, we organize the questions into a chain to enable a well-designed Chain-of-Thought (Wei et al., 2022; Chu et al., 2023): First, we ask questions about the map environment, the ego and other agents’ motion. These questions would let the GPT provide Q&As useful for fine-tuning scene description agents while enhancing the GPT’s understanding of the input information at the same time. Then, we ask questions about the interactions that occur in the observable future period. To include traffic rule-induced and intention-induced interactions, we ask the GPT to think about the interaction between each pair of agents by answering a sequence of questions: Q1: Are the two agents vehicles, and are they close to each other with traffic rules or signals governing their movements? If the answer is yes, we request that GPT analyze their yielding relations according to the traffic rules. Otherwise, we ask Q2: Are the two agents vehicle and pedestrian respectively with intention conflicts? If so, the vehicle generally should yield to the pedestrian for safety. If both questions do not help find the interaction relations, we come to Q3: Does the scene show patterns of certain intention-induced interaction? We provide descriptions of a few common interactions like overtaking, following, etc. GPT can choose the best fit from these patterns to try to cover more interaction types. Based on this pipeline, we can provide interaction analysis with rich information on traffic rule-induced and human intention-induced interactions. After that, we can finally wrap up the interactions between the ego and all other agents to provide the intention of the ego agent.
@@ A.2 Examples of Simulated Vision Modality
Figure 8 provides some whole examples of the simulated scenarios generated by MetaDrive (Li et al., 2022) simulator using ScenarioNet (Li et al., 2023) traffic scenario simulation platform.
A.3 Processing Agent IDs for Motion-LLaVA
Careful processing of agent IDs is essential for enabling Motion-LLaVA to accurately distinguish between different agents. In our approach, these IDs are used purely as textual identifiers within the prompts, allowing the LLM to differentiate agents through natural language understanding. The prompt format is:
[…]
, where <motion> denotes the encoded motion data from each agent’s perspective. The IDs are assigned by range: #0–#99 for vehicles, #100–#199 for bicycles, and #200–#299 for pedestrians. We remap WOMD agent IDs to these local ranges to prevent downstream models from overfitting to specific agents’ behaviors. Importantly, these agent IDs do not encode any additional information beyond their role as identifiers.
A.4 Hyperparameters for Multipath++
In the experiment using languages from Motion-LLaVA outputs to assist vehicle trajectory predictions, we use a batch size of $128$ and a learning rate of $41e-$ 4 for Multipath++. As language is not available for every agent, we use "No captions available for this agent." for those agents without language inputs. The $t5-v1\_1-xxl$ model is used as the language encoder.
A.5 Prompts for evaluating GPT Score
We employ the OpenAI GPT-4-turbo model (OpenAI et al., 2024) to evaluate the performance of all models benchmarked on the WOMD-Reasoning. The GPT Score prompt is presented in Table 12. Notably, we design our GPT Score prompt using a similar template to the one from (Chen et al., 2023).
@@ A.5 Prompts for evaluating GPT Score
A.6 Qualitative Comparison of Benchmarked Multi-modal Models on Answering Questions in WOMD-Reasoning
@@
A.7 Additional Baseline Language Models on WOMD-Reasoning
Our main results in Table 4 primarily evaluate different variants of LLaVA (Liu et al., 2024) to demonstrate the effectiveness of WOMD-Reasoning. For clarity, we explain the categories of questions assessed. The questions are grouped into two main types: factual questions and interaction reasoning questions.
  • Factual questions:
    • Environment-related: Questions about the scenario environment, such as the existence and type of intersections, the number of lanes on the ego vehicle’s side, and the presence or location of crosswalks and stop signs.
    • Ego agent-related: Questions regarding the ego vehicle’s properties, including speed, motion status, lane position, and direction.
    • Surrounding agents-related: Questions about other agents in the scenario, covering their type, speed, motion status, and position relative to the ego vehicle and the intersection.
  • Interaction reasoning questions:
    • Questions about the interactions that will occur between each surrounding agent and the ego agent.
    • Questions about the overall intention of the ego agent in the given scenario.
We further evaluate additional language models (LMs) to confirm the effectiveness of WOMD-Reasoning. Specifically, we assess the answer quality of several baseline models, including LLaMA-Adapter v2.1 (Gao et al., 2023), VITA-1.5 (Fu et al., 2025), and Qwen2.5-VL (Bai et al., 2025) (all 7B versions), both before and after fine-tuning on WOMD-Reasoning. The results are summarized in Table 13.
We observe that without fine-tuning, all models can hardly answer driving-related questions, supporting the motivation of building WOMD-Reasoning. We then fine-tune LLaMA-Adapter on WOMD-Reasoning, which also significantly benefit from the fine-tuning. Besides, we find that fine-tuned our Motion-LLaVA works better than fine-tuned LLaVA or LLaMA-adapter, proving its well-designed structure in utilizing WOMD-Reasoning information.
Table 13: Additional language model baselines evaluated on WOMD-Reasoning. We report the answer quality of several vanilla and WOMD-Reasoning-fine-tuned models, including LLaMA-Adapter, ViTA, and Qwen2.5-VL (all 7B versions), to further validate the effectiveness of WOMD-Reasoning.
Model Fine-tuned on WOMD-Reasoning ROUGE ($\uparrow$) BLEU ($\uparrow$) METEOR ($\uparrow$) CIDEr ($\uparrow$) SPICE ($\uparrow$) GPT Score ($\uparrow$)
LLaVA 0.512 0.211 0.275 1.36 0.455 2.31
LLaMA-Adapter v2.1 0.413 0.174 0.235 0.91 0.372 1.62
ViTA-1.5 0.278 0.121 0.190 0.40 0.227 1.77
Qwen2.5-VL 0.384 0.156 0.247 0.62 0.379 2.36
LLaVA 0.779 0.581 0.439 5.51 0.735 6.88
LLaMA-Adapter v2.1 0.722 0.470 0.375 4.72 0.691 5.14
Motion-LLaVA (Ours) 0.792 0.616 0.449 5.69 0.744 7.02
A.8 Example Aggregated Context for Chain-of-Thought Reasoning
@@
A.9 Example Factual Description Context for Chain-of-Thought Reasoning
@@
A.10 Ablations on Motion-LLaVA
Chain-of-Thought Inference. To address the hallucinated response problem (Gai et al., 2024; Tonmoy et al., 2024; Sima et al., 2023) in reasoning interactions, we use a Chain-of-Thought (CoT) approach to provide Motion-LLaVA with structured context for more accurate reasoning. Table 14 compares model performance with and without CoT on interaction and intention analysis. We see that with CoT, the model performs better in GPT Scores. We further show that if our motion data translation program A.1.1 is used during inference to provide contexts, i.e., with factual description context (an example can be found in Appendix A.9) and CoT, the performance can be significantly improved, proving the potential of the CoT inference strategy.
Table 14: Effects of CoT in answering reasoning-related questions.
Direct 0.616 0.474 0.365 2.54 0.578 6.59
CoT 0.614 0.474 0.366 2.52 0.571 6.76
CoT w/ Facts Contexts 0.664 0.542 0.403 3.10 0.631 7.64
Influence of Pre-training for the Motion Vector Encoder. To establish a pre-training strategy for the motion vector encoder, we examine factual question answering across Motion-LLaVA with four pre-training configurations: no pre-training, autoencoder-style pre-training, motion-caption pre-training (Chen et al., 2023), and pre-training with a motion prediction task. All configurations shared the same encoder structure. In the motion-caption pre-training, we pre-train all but LLM weights in Motion-LLaVA, by penalizing errors in captioning the motion. For autoencoder-style pre-training and pre-training with a motion prediction task, we retain only the encoder after pre-training, to function as the motion vector encoder for Motion-LLaVA.
None 0.814 0.684 0.491 5.95 0.774 30.0% 2.00 6.79
Motion Caption 0.824 0.711 0.498 6.05 0.775 42.6% 2.24 6.97
AutoEncoder 0.837 0.727 0.516 6.35 0.803 47.3% 1.41 7.07
Motion Prediction (Ours) 0.840 0.736 0.516 6.35 0.794 56.2% 1.00 7.09
Motion-LLaVA Fine-tuning Strategy. The original LLaVA (Liu et al., 2024) training pipeline consists of two stages: a projector alignment stage (also known as pre-training) to align vision and language features, followed by a fine-tuning stage to enhance instruction-following capabilities. Throughout this process, the vision encoder remains frozen. Since our Motion-LLaVA builds on the architecture of the original LLaVA (Liu et al., 2024), it seems reasonable to adopt a similar training approach for the WOMD-Reasoning task.
However, based on extensive experiments with fine-tuning strategies in Table 16, we find that incorporating projector alignment stage or keeping the motion vector encoder frozen throughout, which are practices employed in LLaVA (Liu et al., 2024), resulted in degraded performance on WOMD-Reasoning. Consequently, our Motion-LLaVA adopts a single-stage fine-tuning approach, where the motion vector encoder is unfrozen.
Table 16: Refinement of Motion-LLaVA training strategy.
LLaVA 0.821 0.678 0.484 5.82 0.771 2.3% 12.00 6.17
$\bm{-}$ Projector Alignment 0.821 0.697 0.492 5.96 0.782 2.3% 8.06 6.49
A.11 WOMD-Reasoning Dataset Efficiency Analysis
To assess the data efficiency of WOMD-Reasoning, we conduct scaling studies to analyze how model performance improves as the used dataset’s size increases, from 0.5M to 1M and up to 3M Q&A pairs (full dataset).
In Table 17, we find that with more WOMD-Reasoning data involved, the fine-tuned model first gains the ability to answer factual questions, which hits a good GPT score with only 0.5M data. As the dataset grows, the model gradually learns to answer interaction reasoning questions. These justify the size of WOMD-Reasoning, as well as its data efficiency in helping models to answer hierarchical driving-related questions.
Table 17: Performance of Motion-LLaVA with different WOMD-Reasoning dataset sizes on factual and interaction reasoning questions. Results confirm the data efficiency of WOMD-Reasoning in improving driving-related reasoning.
Dataset Size Factual Questions Interaction Reasoning Questions
ROUGE BLEU METEOR CIDEr SPICE GPT ROUGE BLEU METEOR CIDEr SPICE GPT
0.5M 0.806 0.683 0.477 5.77 0.760 6.69 0.562 0.410 0.339 1.94 0.513 5.67
1M 0.818 0.701 0.491 5.90 0.773 6.74 0.589 0.447 0.354 2.23 0.547 6.36
3M 0.840 0.736 0.516 6.35 0.794 7.09 0.614 0.474 0.366 2.52 0.571 6.76